OpenStudy (anonymous):

second derivtive of xy+y^2=1. got a complicated ans. please help with this.

OpenStudy (amistre64):

xy+y^2=1 x'y + xy' +2y y' = 0 x''y + x'y' + x'y' + xy'' +2 y' + 2y y'' = 0

OpenStudy (anonymous):

(xy+2y+x)/(2y+3)^3 i got this as 2nd derivative

OpenStudy (amistre64):

since x' = 1; and x'' = 0 y' + y' + xy'' +2 y' + 2y y'' = 0 4y' + y''(x + 2y) = 0 y'' = -4y'/(x + 2y)

OpenStudy (ash2326):

need to find d^2y/dx^2 first find dy/dx xy'+y+2yy'=0 now differentiate again 2y'+xy''+2y'^2+2yy''=0 we get y''(x+2y)=-(2y'+2y'^2) y''=-(2y'+2y'^2)/(x+2y)

OpenStudy (amistre64):

dealing with the y' x'y + xy' +2y y' = 0 y + y'(x +2y) = 0 y' = -y/(x +2y) .......................... y'' = -4y'/(x + 2y) y'' = -4(-y/x+2y)/(x + 2y) y'' = 4y/(x + 2y)^3 is what i get, but id have to dbl check

OpenStudy (anonymous):

1st derivative is the same as yours

OpenStudy (anonymous):

\[y+xy'+2yy'=0\] \[(x+2y)y'=-y\] \[y'=\frac{y}{2x+y}\] is that i get, then one more time

OpenStudy (anonymous):

oops i am off by a minus sign at the end

OpenStudy (anonymous):

for2nd derivative i used quotient rule

OpenStudy (anonymous):

yes, and then you will have y' in your answer, so you should substitute back \[y'=\frac{-y}{x+2y}\] and do some more algebra. big fun

OpenStudy (anonymous):

no fun at all. i think i messed up in algebra

OpenStudy (anonymous):

it looks like amistre did it already. i am late as usual

OpenStudy (anonymous):

ugh

OpenStudy (amistre64):

:) but i reserve the right to be wrong lol

OpenStudy (anonymous):

ok i get \[\frac{y-xy'}{(x+2y)^2}\] as a first step. look good?

OpenStudy (anonymous):

now i'm getting (xy+y^2+y^2+xy)/(2y+3)^3. replacing the xy+y^2 with 1 frm question. 2/(2y+3)^3

OpenStudy (anonymous):

is that ok

OpenStudy (anonymous):

could be but it is not what i got. i am fairly sure of this line \[\frac{y-xy'}{(x+2y)^2}\]

OpenStudy (amistre64):

mines off a bit ... bummer

OpenStudy (anonymous):

oh once again i am off by a minus sign. let me check again

OpenStudy (anonymous):

ok i got what i think is right \[\frac{y+\frac{xy}{x+2y}}{(x+2y)^2}=\frac{y(x+2y)+xy}{(x+2y)^3}=\frac{xy+2y^2+xy}{(x+2y)^3}=\frac{1+xy}{(x+2y)^3}\]

OpenStudy (anonymous):

yep that's what i got as well. thanx